OpenStudy (anonymous):

How can (f(3+h)-f(3))/h equal both 1 and -1? How do I solve for this equation I forgot how to do this!! Please help huge exam is tomorrow!!

OpenStudy (jack1):

im missing something here dude... (f(3+h)-f(3))/h \[\frac {f(3+h)-f(3)}{h}\] is this one of those equations of (as h approaches zero?)

OpenStudy (anonymous):

yeah sorry dude

OpenStudy (anonymous):

it says as h-->0+ it's 1 and as h-->0- it's -1. How do you find that out and how do you even use algebra to solve this?

OpenStudy (jack1):

do you have the equation for f(x)?

OpenStudy (anonymous):

no? isn't that f(x)?

OpenStudy (jack1):

f(x) is any equation (f(x) is just * a function of x is...* ie f(x) = x^2 + 37 so f(3) = 3^2 + 37 = 9 + 37

OpenStudy (jack1):

i think we may need the equation for f(x) to solve..., hang on I'll ask @satellite73 ...dude u have a sec?

OpenStudy (anonymous):

i think it's just a general question, you don't need an equation. I think what they mean to say is for any f(x) it's possible for that to equal 1 and -1

OpenStudy (jack1):

ok... but assuming : f(x) = x^2 + 37 at x = 3 y = (f(x+h)-f(x))/h =(( (x+h)^2 + 37)- (x^2 + 37))/h \[\frac {f(x+h)-f(x)}{h}\] \[\frac {((x+h)^2 + 37) - (x)^2 + 37}{h}\] \[\frac {(x^2+2xh+h^2 + 37) - (x)^2 + 37}{h}\] when x = 3 \[\frac {(3^2+2*3*h+h^2 + 37) - (3)^2 + 37}{h}\] \[\frac {9+6h+h^2 + 37 - 9 - 37}{h}\] \[\frac {9+6h+h^2 + 37 - 9 - 37}{h}\] \[\frac {6h+h^2}{h}\] \[6+h\] so as h --->0 y = 6 i remember now, this is a method for determining a derivative

OpenStudy (jack1):

so if your f(x) is x^2 + 37 f'(x) = 2x so f'(3) = 2*3 = 6

OpenStudy (anonymous):

exactly I don't get it wtf

OpenStudy (jack1):

have you covered derivatives before in class?

OpenStudy (jack1):

all its saying is that at the point x = 3, the slope (gradient) of your equation is either +1 or -1 so your equation is probably something like f(x) = 1/x or 1/sqrt x or something similar

OpenStudy (jack1):